OpenStudy (anonymous):

a top-class footballer earns £7.94 per minute calculate the amount this footballer earns per year (365 days), give ur answer in ordinary notation unrounded, then in scientific notation correct to two significant figures

OpenStudy (anonymous):

Well, there are of course 1440 minutes every day, so, we could write this as an equation, 1440 minutes, times 365 days, times 7.94.

OpenStudy (anonymous):

\(\frac{7.94}{min} \frac{60 min}{hr} \frac{24 hr}{day} \frac{365 day}{year}\)

OpenStudy (anonymous):

need help still?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

of course, ok, tell me.

OpenStudy (anonymous):

oh. Sorry, I do not know what scientific notation is, so that was kinda useless.

OpenStudy (anonymous):

I'm afraid I cant help you.

OpenStudy (anonymous):

how can i put it in a scientific notation?

OpenStudy (anonymous):

Ill see what I can do, so, is scientific notation is like what the man wrote?

OpenStudy (anonymous):

ok thanks guys

OpenStudy (anonymous):

oh, I see now

OpenStudy (anonymous):

\(5.12 \times 10^6\)

OpenStudy (anonymous):

thanks mark

OpenStudy (anonymous):

I should have read more closely, two significant figures would be \(5.1 \times 10^6\)

OpenStudy (anonymous):

i got that, how can u write it in significant figures?

OpenStudy (anonymous):

\(5.119 \times 10^6\) is four significant figures \(5.12 \times 10^6\) is three significant figures \(5.1 \times 10^6\) is two significant figures \(5 \times 10^6\) is one significant figures

OpenStudy (anonymous):

if there are 25000 people watching the football match,how much would each person pay to cover the player's salary?